A hydroelectric power plant uses spinning turbines to transform the kinetic energy of moving water into
electric energy with 80 % efficiency. That is, 80 % of the kinetic energy becomes electric energy. A small
hydroelectric plant at the base of a dam generates 50 MW of electric power when the falling water has a
speed of 18 m/s.
What is the water flow rate -kilograms of water per second-through turbines?
1
Expert's answer
2020-05-07T09:49:28-0400

Solution.

"P=50\sdot10^6W;"

"\upsilon=18m/s;"

"\eta=80" %;

"P=\dfrac{W}{t};\implies W=Pt;"

Since 80% of kinetic energy becomes electrical, we have:

"\eta=\dfrac{W}{E_k}\sdot100" %; "\implies E_k=\dfrac{W}{\eta}\sdot100" %;

Kinetic energy is calculated by the formula:

"E_k=\dfrac{m\upsilon^2}{2}; \dfrac{m\upsilon^2}{2}=\dfrac{Pt}{ \eta}\sdot100" %;"\implies"

Water consumption per second is equal to:

"\dfrac{m}{t}=\dfrac{2P}{\eta\upsilon^2}\sdot100" %;

"\dfrac{m}{t}=\dfrac{2\sdot50\sdot10^6W}{80\%\sdot(18m/s)^2}\sdot100\%=3.86\sdot10^5kg/s;"

Answer:"\dfrac{m}{t}=3.86\sdot10^5kg/s;"